I have to use the distributive property to simplify these problems, please let me know if these are correct.
98*7= 7(90+8)
9*28 = 9(20+8)
78*8 = (70+8)8
7(2,009) = 7(2,009)
899*5 = (800 +99)5
30*105 = 30(100*5)
8*5-12*5 = 5(5+3)-5(10+2) or 5(8)-5(12)
7*10+7(-3) = 7(10+3)
-4(3)+(-4)(6) = -4(3+6)
6(8) +6(-2) = 6(8+2)

Found 3 solutions by consc198, ikleyn, timofer:
Answer by consc198(59)

Yes they are correct

Answer by ikleyn(53996)

.
I have to use the distributive property to simplify these problems, please let me know if these are correct.
98*7= 7(90+8)
9*28 = 9(20+8)
78*8 = (70+8)8
7(2,009) = 7(2,009)
899*5 = (800 +99)5
30*105 = 30(100*5)
8*5-12*5 = 5(5+3)-5(10+2) or 5(8)-5(12)
7*10+7(-3) = 7(10+3)
-4(3)+(-4)(6) = -4(3+6)
6(8) +6(-2) = 6(8+2)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The other person (@consc198) says in his post  "Yes they are correct".

        Be aware:  his answer is  WRONG.

        See my solution below with my comments.


98*7= 7(90+8)                            correct.

9*28 = 9(20+8)                           correct.

78*8 = (70+8)8                           correct.

7(2,009) = 7(2,009)                      Probably, it is 7*2.009 = 7*2 + 7*0.009.  In this form, it is correct.

899*5 = (800 +99)5                       I would simplify it THIS WAY:  899*5 = (900-1)*5 = 900*5 - 5.

30*105 = 30(100*5)                       In this form, it is wrong.  The correct form is  30*105 = 30(100+5).

8*5-12*5 = 5(5+3)-5(10+2) or 5(8)-5(12)  My interpretation is 8*5 - 12*5 = (8-12)*5.

7*10+7(-3) = 7(10+3)                     Wrong. A correct form is 7*10 + 7*(-3) = 7*(10-3).

-4(3)+(-4)(6) = -4(3+6)                  correct.

6(8) +6(-2) = 6(8+2)                     Wrong.  A correct form is  6(8) + 6(-2) = 6(8-2).

Solved correctly.

As I read and checked several solutions by @consc198, I see that this person produces wrong solutions without any hesitation.

By producing gibberish, he invites readers to visit his web-site for more help.

By observing the quality of his solutions, I would stay a cannon shot away from his website.



Answer by timofer(159)

98*7= 7(90+8)
wrong

9*28 = 9(20+8)
wrong

78*8 = (70+8)8
wrong

7(2,009) = 7(2,009)
wrong

899*5 = (800 +99)5
wrong

30*105 = 30(100*5)
wrong

8*5-12*5 = 5(5+3)-5(10+2) or 5(8)-5(12)
strange or wrong

7*10+7(-3) = 7(10+3)
wrong

-4(3)+(-4)(6) = -4(3+6)
CORRECT

6(8) +6(-2) = 6(8+2)
wrong